# Group Projected Subspace Pursuit for Block Sparse Signal Reconstruction: Convergence Analysis and Applications

### v1

@@ 3 GPSP and Main Results
The Group Projected Subspace Pursuit (GPSP) algorithm was first proposed in [25] and used in [26] for identifying PDEs with varying coefficients. As its algorithmic design is independent from the particular form associated with the PDE identification, GPSP can also be used for general block-sparse signal reconstruction. Algorithm 1 presents the details for GPSP [^1] In [25], the metric for initialization and expansion was defined by $|\text{proj}(\mathbf{v},\mathbf{F}_{g})|/(\|\mathbf{v}\|_{2}\|\text{proj}(\mathbf{v},\mathbf{F}_{g})\|_{2})$ for $\mathbf{v}=\mathbf{y}$ or $\mathbf{y}_{r}^{l-1}$, which equals $\|\text{proj}(\mathbf{v},\mathbf{F}_{g})\|_{2}/\|\mathbf{v}\|_{2}$. In this paper, we present this equivalent form for convenience.. It shares similar structures with state-of-the-art greedy algorithms, such as BSP [29], and we compare them in Section 4.
In this section, our focus is on proving the convergence and robustness of GPSP, thus theoretically justifying its validity in recovering block-sparse signals for general purposes. In Section 3.1, we present our main result (Theorem 1) stating a sufficient condition for the exact recovery property of GPSP. In Section 3.2, we analyze the behaviors of GPSP when the observation vector is perturbed by noise. We outline the major steps for proving Theorem 1 and Theorem 2 in Section 3.3 and Section 3.4, respectively, and delegate the details in the Appendix. We note that the existing proof for BSP in [29] cannot be trivially applied to analyze GPSP as different criteria are used for selecting and removing features; new techniques need to be developed. In the following, we denote the set of group indices selected by GPSP as $\widehat{T}$. The corresponding reconstruction $\widehat{\mathbf{c}}$ is obtained by setting entries associated with blocks indexed by $\widehat{T}$ via
@@
Algorithm 1 Group Projected Subspace Pursuit (GPSP) [25]
@@ 3.1 Convergence of GPSP
Theorem 1 is proved in Section 3.3. Theorem 1 provides a sufficient condition for the convergence of GPSP in terms of the BRIC $\delta_{M,2k}$ of the block matrix $\mathbf{A}$. In particular, for block matrices with sufficiently small BRIC $\delta_{M,2k}$, GPSP is guaranteed to converge. Figure 2 (a) plots the graph of $C_{M,k}$ in (3.2) against $\delta_{M,2k}$; it increases as $\delta_{M,2k}$ gets larger before it reaches the value of $1$. Using the bisection method, we deduce that GPSP converges when $\delta_{M,2k}$ is bounded by approximately $0.1188$. To guarantee an exact recovery of a block $k$-sparse signal, we note that the BRIC in (3.2) is related to the block sparsity level $2k$. This is consistent with the analysis of the uniqueness of block sparse representations. See Appendix A for details. We also remark that a similar sufficient condition is available for BSP [29], which is guaranteed to converge when $\delta_{M,3k}<0.1672$. Although the upper-bound constant is higher than $0.1188$, their condition is associated with the BRIC of order $3k$, and the relation $\delta_{M,2k}\leq\delta_{M,3k}$ makes it difficult to judge which condition is more relaxed. Instead, we shall compare GPSP with BSP and other related algorithms in Section 4 and Section 5 via various examples and experiments.
@@ 3.2 Recovery distortion with perturbed measurements
Theorem 2 is proved in Section 3.4. Theorem 2 is a generalization of the result in [10] (Theorem 9) for GPSP. It states that the distortion of the recovery is bounded by the magnitude of the observational perturbation multiplied by a factor. As shown in Figure 2 (b), the factor $G_{M,k}$ in (3.3) decreases as $\delta_{M,2k}$ increases. When $\delta_{M,2k}$ approaches the bound $0.0916$, $G_{M,k}$ is approximately $13.9825$; when $\delta_{M,2k}$ approaches 00 from right, $G_{M,k}$ grows to infinity. This indicates that mild correlations among blocks can be helpful for handling noise. Theorem 2 justifies the robustness of GPSP against noise for sampling matrices with sufficiently small BRIC, and the reconstruction error reduces to 00 when the perturbation vanishes. In particular, the relative $L_{2}$-error $\|\mathbf{c}^{*}-\widehat{\mathbf{c}}\|_{2}/\|\mathbf{c}^{*}\|_{2}<\alpha\%$ whenever $\|\mathbf{e}\|_{2}<13.9825^{-1}\times\alpha\%\|\mathbf{c}^{*}\|_{2}=0.0715\times\alpha\%\|\mathbf{c}^{*}\|_{2}$; hence, GPSP has a higher tolerance for the noise if the true signal $\mathbf{c}^{*}$ has a larger magnitude than that of the noise.
@@ 3.3 Outline of the proof for Theorem 1
Our strategy of proving the convergence of GPSP is inspired by [10, 29]. As stated in Algorithm 1, each iteration of GPSP contains two stages: a dictionary expanding stage (line 5) and a dictionary shrinking stage (line 7). Hence, there are three key criteria towards the establishment of Theorem 1: (i) correct features need to be included during the expansion (Theorem 3); (ii) correct features need to survive during the shrinking (Theorem 4); and (iii) the fitting errors need to reduce (Theorem 5).
@@ 3.4 Outline of the proof for Theorem 2
The proof for Theorem 2 is analogous to [10, Proof for Theorem 2]. With the measurement perturbation, we shall first bound the recovery distortion as follows.
@@ 4 Relations to Other Greedy Algorithms
BOMP [14] $\varnothing$ $\arg\max_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$
BOMPR [18] $\varnothing$ $\arg\min_{g}\|\mathbf{y}_{r}^{l-1}-\text{proj}(\mathbf{y}_{r}^{l-1},\mathbf{F}_{g})\|_{2}$
BCoSaMP [52] $\varnothing$ $\arg\text{top}^{2k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{x}_{p}^{l}[g]\|_{2}$
BSP [29] $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{x}_{p}^{l}[g]\|_{2}$
GPSP [25] $\arg\text{top}^{k}_{g}\|\text{proj}(\mathbf{y},\mathbf{F}_{g})\|_{2}$ $\arg\text{top}^{k}_{g}\|\text{proj}(\mathbf{y}_{r}^{l-1},\mathbf{F}_{g})\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}\mathbf{x}_{p}^{l}[g]\|_{2}$
@@
In the initialization stage, some algorithms may start with none, e.g., BOMP [14], BOMPR [18], and BCoSaMP [52], and others have initial guesses, e.g., BSP [29] and GPSP [25]. During the expansion stage, they select candidates according to different criteria. Algorithms like BOMP and BOMPR select one candidate at a time, whereas BCoSaMP, BSP, and GPSP select multiple candidates. During the shrinking stage, BOMP and BOMPR do not exclude chosen candidates; BCoSaMP, BSP, and GPSP refine their selections by removing chosen candidates according to certain rules. Table 1 summarizes these discrepancies among BOMP, BOMPR, BCoSaMP, BSP, and GPSP for a systematic comparison.
@@
Finally, we note that RCC is a direct extension of the criterion used in CoSaMP [39] and SP [10], while RMC does not follow the same paradigm. We numerically compare these two criteria in Subsection 5.6 via a group of ablation studies.
@@ 5 Numerical Experiments
In this section, we present various experiments to validate the effectiveness of GPSP by conducting a series of comparison and ablation studies under diverse scenarios. We focus on comparing GPSP [25] with BOMP [14], BOMPR [18], BCoSaMP [52], and BSP [29]. We test these methods in practical cases where block features have different statistical properties with and without noise perturbations. We investigate the algorithmic performances on random sampling matrices with non-Gaussian types. Furthermore, we consider more complicated settings in face recognition and PDE identification to test GPSP and state-of-the-art methods. All the algorithms are implemented in Python and available online222 https://github.com/RoyYuchenHe/BlockSparse.
@@ 5.1 Heterogeneous Gaussian blocks
Figure 4 shows the experiment results. Although the probability of exact recovery for each method reduces as the underlying signals have fewer zero blocks, we observe that GPSP outperforms other methods for varying levels of block sparsity and different block sizes. In particular, we find that BOMP has almost less than $50\%$ of success rate for all levels of sparsity, with and without column normalization. Such unsatisfying behaviors of BOMP were also reported in [18] when the blocks are redundant. We note that BOMPR presents better recovery results than BOMP, further validating the effectiveness of using projection for feature selection, as discussed in Section 4. Generally better than BOMP, the success rates of BSP and BCoSaMP are improved by column normalization, although they are consistently lower than those of BOMPR and GPSP. For signals with single non-zero blocks, both GPSP and BOMPR identify the correct features in all experiments with and without column normalization. For lower levels of sparsity ($2\sim 5$), both GPSP and BOMPR keep their success rates above $50\%$. For middle levels of sparsity ($6\sim 8$), we observe that the success rate of BOMPR drops below $50\%$ especially when the columns are not normalized; whereas the success rate of GPSP still maintains above $50\%$. For higher levels of sparsity (above $10$), the success rate of GPSP slowly decreases while remaining above $20\%$, and the success rate of BOMPR stays below that of GPSP in all settings.
@@ 5.4 Application to Face Recognition
One of the applications of block sparse signal recovery in image processing is face recognition [30, 16]. Given a face image, the goal is to identify the person from a list of subjects whose face images under different lighting conditions are archived. For each subject, the corresponding archived face images are grouped as a block feature, and our sampling matrix, or dictionary, comprises all the block features; and we can treat the test image as the response. Under the Lambertian assumption that each image is represented as a linear combination of other images of its subject in the dictionary, the problem of face recognition can be addressed by finding the block, i.e., the subject, whose feature matches the test image, or equivalently, by reconstructing a signal with block sparsity one.
We evaluate and compare the performances of GPSP with other algorithms using the Extended Yale B dataset [32] derived from [20]. This dataset consists of face images belonging to 38 human subjects. Each subject is represented by 64 face images, which are captured under various illuminating conditions. The first row of Figure 2 shows 8 images of a specific subject; they exemplify the variations in the dataset, including contrast change (the 2nd image), minor differences in facial expressions (the 3rd image), and extremely low lighting (the 5th image). For each subject, we randomly sampled $M$ images to form our training dataset of size $38\times M$, and we use the rest images for evaluation. Upon stretching each image to a vector of length $32256\,(=192\times 168)$, we employ three strategies to reduce the dimension to $D=132$ as in [16]:
  • 1.
@@
Table 2: Comparison study on the task of face recoginition using the Extended Yale B dataset [32]. The first row shows sample images of a single subject. The table records the average values of frequency of exact recovery over 10 trials for different training size per subject $M=9,18,25,32$ and three dimension reduction strategies: principle component analysis (PCA), random projection (randomProj), and down-sampling (downSample). The highest levels of accuracy are bolded, and the second best are underlined.
@@ 5.5 Application to PDE Identification
GPSP was first proposed in [24] for identifying PDEs with time-space dependent coefficients from trajectory data. It was shown that GPSP can find the underlying models for various important classes of PDEs with higher success rates than BSP. In this experiment, we extend the comparison by including also BOMP, BOMPR, and BCoSaMP. We consider the viscous Burgers’ equation
@@
with $\phi(t)=0.5+0.5\tanh(10(t-0.15))$. Following the numerical settings in [25], we discretize the computational domain with spatial step 0.02 and temporal step $3\times 10^{-4}$ with which the PDE is solved. Then we downsample the data to spatial step 0.04 and temporal step $3\times 10^{-3}$, such that the data is over a $50\times 100$ grid.
The PDE (5.1) has non-trivial behaviors. The coefficient $a(x,t)$ is visualized in Figure 7 (a), and the corresponding solution obtained by the spectral method (See [24]) is shown in (b). It models a non-linear advection with a homogeneous diffusion on a circle. The speed of the advection is affected by the solution itself as well as the varying coefficient $a$. When the solution $u$ is positive, it moves from right to left; when it is negative, it moves from left to right. Moreover, the higher the magnitude of the solution is, the faster it moves locally. Consequently, solutions can form shocks without diffusion. For more details about the properties of the PDE, we refer the readers to [17].
We set up a dictionary containing candidate differential operators as features to identify the PDE from the solution data. We include partial derivatives of $u$ up to order 4 and products of no more than 3 terms, leading to 55 features in the dictionary. Since the coefficients may vary both in space and time, we approximate them using linear combinations of B-spline bases of order $2$. As a result, each feature is represented by a block matrix whose columns are associated with local bases. We use 7 bases for space and 8 bases for time, leading to a block size of 56. The details of the implementation are explained in [24]. Assuming that we know the sparsity is 2, we apply BOMP, BOMPR, BCoSaMP, BSP, and GPSP to find the features. Table 3 shows the identified features by these methods. In this experiment, only GPSP finds the correct features, giving the underlying viscous Burgers’ equation. BOMP and BOMPR have identical results; they find the correct second-order differential operator, yet the non-linear part is wrong. BCoSaMP and BSP fail to identify any correct features, and the respective models are more complicated than the true equation. This comparison verifies the effectiveness of GPSP in the application of PDE identification.

### v2

@@ 3 GPSP and Main Results
The Group Projected Subspace Pursuit (GPSP) algorithm was first proposed in [HKL+23] and used in [HZZ23] for identifying PDEs with varying coefficients. As its algorithmic design is independent from the particular form associated with the PDE identification, GPSP can also be used for general block-sparse signal reconstruction. Algorithm 1 presents the details for GPSP [^1] In [HKL+23], the metric for initialization and expansion was defined by $|\text{proj}(\mathbf{v},\mathbf{F}_{g})|/(\|\mathbf{v}\|_{2}\|\text{proj}(\mathbf{v},\mathbf{F}_{g})\|_{2})$ for $\mathbf{v}=\mathbf{y}$ or $\mathbf{y}_{r}^{l-1}$, which equals $\|\text{proj}(\mathbf{v},\mathbf{F}_{g})\|_{2}/\|\mathbf{v}\|_{2}$. In this paper, we present this equivalent form for convenience.. It shares similar structures with state-of-the-art greedy algorithms, such as BSP [KSHT13], and we compare them in Section 4.
In this section, our focus is on proving the convergence and robustness of GPSP, thus theoretically justifying its validity in recovering block-sparse signals for general purposes. In Section 3.1, we present our main result (Theorem 1) stating a sufficient condition for the exact recovery property of GPSP. In Section 3.2, we analyze the behaviors of GPSP when the observation vector is perturbed by noise. We outline the major steps for proving Theorem 1 and Theorem 2 in Section 3.3 and Section 3.4, respectively, and delegate the details in the Appendix. We note that the existing proof for BSP in [KSHT13] cannot be trivially applied to analyze GPSP as different criteria are used for selecting and removing features; new techniques need to be developed. In the following, we denote the set of group indices selected by GPSP as $\widehat{T}$. The corresponding reconstruction $\widehat{\mathbf{c}}$ is obtained by setting entries associated with blocks indexed by $\widehat{T}$ via
@@
Algorithm 1 Group Projected Subspace Pursuit (GPSP) [HKL+23]
@@ 3.1 Convergence of GPSP
Theorem 1 is proved in Section 3.3. Theorem 1 provides a sufficient condition for the convergence of GPSP in terms of the BRIC $\delta_{M,2k}$ of the block matrix $\mathbf{A}$. In particular, for block matrices with sufficiently small BRIC $\delta_{M,2k}$, GPSP is guaranteed to converge. Figure 2 (a) plots the graph of $C_{M,k}$ in (3.2) against $\delta_{M,2k}$; it increases as $\delta_{M,2k}$ gets larger before it reaches the value of $1$. Using the bisection method, we deduce that GPSP converges when $\delta_{M,2k}$ is bounded by approximately $0.1188$. To guarantee an exact recovery of a block $k$-sparse signal, we note that the BRIC in (3.2) is related to the block sparsity level $2k$. This is consistent with the analysis of the uniqueness of block sparse representations. See Appendix A for details. We also remark that a similar sufficient condition is available for BSP [KSHT13], which is guaranteed to converge when $\delta_{M,3k}<0.1672$. Although the upper-bound constant is higher than $0.1188$, their condition is associated with the BRIC of order $3k$, and the relation $\delta_{M,2k}\leq\delta_{M,3k}$ makes it difficult to judge which condition is more relaxed. Instead, we shall compare GPSP with BSP and other related algorithms in Section 4 and Section 5 via various examples and experiments.
@@ 3.2 Recovery distortion with perturbed measurements
Theorem 2 is proved in Section 3.4. Theorem 2 is a generalization of the result in [DM09] (Theorem 9) for GPSP. It states that the distortion of the recovery is bounded by the magnitude of the observational perturbation multiplied by a factor. As shown in Figure 2 (b), the factor $G_{M,k}$ in (3.3) decreases as $\delta_{M,2k}$ increases. When $\delta_{M,2k}$ approaches the bound $0.0916$, $G_{M,k}$ is approximately $13.9825$; when $\delta_{M,2k}$ approaches 00 from right, $G_{M,k}$ grows to infinity. This indicates that mild correlations among blocks can be helpful for handling noise. Theorem 2 justifies the robustness of GPSP against noise for sampling matrices with sufficiently small BRIC, and the reconstruction error reduces to 00 when the perturbation vanishes. In particular, the relative $L_{2}$-error $\|\mathbf{c}^{*}-\widehat{\mathbf{c}}\|_{2}/\|\mathbf{c}^{*}\|_{2}<\alpha\%$ whenever $\|\mathbf{e}\|_{2}<13.9825^{-1}\times\alpha\%\|\mathbf{c}^{*}\|_{2}=0.0715\times\alpha\%\|\mathbf{c}^{*}\|_{2}$; hence, GPSP has a higher tolerance for the noise if the true signal $\mathbf{c}^{*}$ has a larger magnitude than that of the noise.
@@ 3.3 Outline of the proof for Theorem 1
Our strategy of proving the convergence of GPSP is inspired by [DM09, KSHT13]. As stated in Algorithm 1, each iteration of GPSP contains two stages: a dictionary expanding stage (line 5) and a dictionary shrinking stage (line 7). Hence, there are three key criteria towards the establishment of Theorem 1: (i) correct features need to be included during the expansion (Theorem 3); (ii) correct features need to survive during the shrinking (Theorem 4); and (iii) the fitting errors need to reduce (Theorem 5).
@@ 3.4 Outline of the proof for Theorem 2
The proof for Theorem 2 is analogous to [DM09, Proof for Theorem 2]. With the measurement perturbation, we shall first bound the recovery distortion as follows.
@@ 4 Relations to Other Greedy Algorithms
BOMP [EKB10] $\varnothing$ $\arg\max_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$
BOMPR [FLZZ14] $\varnothing$ $\arg\min_{g}\|\mathbf{y}_{r}^{l-1}-\text{proj}(\mathbf{y}_{r}^{l-1},\mathbf{F}_{g})\|_{2}$
BCoSaMP [ZXC+19] $\varnothing$ $\arg\text{top}^{2k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{x}_{p}^{l}[g]\|_{2}$
BSP [KSHT13] $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}^{\top}\mathbf{y}_{r}^{l-1}\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{x}_{p}^{l}[g]\|_{2}$
GPSP [HKL+23] $\arg\text{top}^{k}_{g}\|\text{proj}(\mathbf{y},\mathbf{F}_{g})\|_{2}$ $\arg\text{top}^{k}_{g}\|\text{proj}(\mathbf{y}_{r}^{l-1},\mathbf{F}_{g})\|_{2}$ $\arg\text{top}^{k}_{g}\|\mathbf{F}_{g}\mathbf{x}_{p}^{l}[g]\|_{2}$
@@
In the initialization stage, some algorithms may start with none, e.g., BOMP [EKB10], BOMPR [FLZZ14], and BCoSaMP [ZXC+19], and others have initial guesses, e.g., BSP [KSHT13] and GPSP [HKL+23]. During the expansion stage, they select candidates according to different criteria. Algorithms like BOMP and BOMPR select one candidate at a time, whereas BCoSaMP, BSP, and GPSP select multiple candidates. During the shrinking stage, BOMP and BOMPR do not exclude chosen candidates; BCoSaMP, BSP, and GPSP refine their selections by removing chosen candidates according to certain rules. Table 1 summarizes these discrepancies among BOMP, BOMPR, BCoSaMP, BSP, and GPSP for a systematic comparison.
@@
Finally, we note that RCC is a direct extension of the criterion used in CoSaMP [NT09] and SP [DM09], while RMC does not follow the same paradigm. We numerically compare these two criteria in Subsection 5.6 via a group of ablation studies.
@@ 5 Numerical Experiments
In this section, we present various experiments to validate the effectiveness of GPSP by conducting a series of comparison and ablation studies under diverse scenarios. We focus on comparing GPSP [HKL+23] with BOMP [EKB10], BOMPR [FLZZ14], BCoSaMP [ZXC+19], and BSP [KSHT13]. We test these methods in practical cases where block features have different statistical properties with and without noise perturbations. We investigate the algorithmic performances on random sampling matrices with non-Gaussian types. Furthermore, we consider more complicated settings in face recognition and PDE identification to test GPSP and state-of-the-art methods. All the algorithms are implemented in Python and available online222 https://github.com/RoyYuchenHe/BlockSparse.
@@ 5.1 Heterogeneous Gaussian blocks
Figure 4 shows the experiment results. Although the probability of exact recovery for each method reduces as the underlying signals have fewer zero blocks, we observe that GPSP outperforms other methods for varying levels of block sparsity and different block sizes. In particular, we find that BOMP has almost less than $50\%$ of success rate for all levels of sparsity, with and without column normalization. Such unsatisfying behaviors of BOMP were also reported in [FLZZ14] when the blocks are redundant. We note that BOMPR presents better recovery results than BOMP, further validating the effectiveness of using projection for feature selection, as discussed in Section 4. Generally better than BOMP, the success rates of BSP and BCoSaMP are improved by column normalization, although they are consistently lower than those of BOMPR and GPSP. For signals with single non-zero blocks, both GPSP and BOMPR identify the correct features in all experiments with and without column normalization. For lower levels of sparsity ($2\sim 5$), both GPSP and BOMPR keep their success rates above $50\%$. For middle levels of sparsity ($6\sim 8$), we observe that the success rate of BOMPR drops below $50\%$ especially when the columns are not normalized; whereas the success rate of GPSP still maintains above $50\%$. For higher levels of sparsity (above $10$), the success rate of GPSP slowly decreases while remaining above $20\%$, and the success rate of BOMPR stays below that of GPSP in all settings.
@@ 5.4 Application to Face Recognition
One of the applications of block sparse signal recovery in image processing is face recognition [LCW17, EV12]. Given a face image, the goal is to identify the person from a list of subjects whose face images under different lighting conditions are archived. For each subject, the corresponding archived face images are grouped as a block feature, and our sampling matrix, or dictionary, comprises all the block features; and we can treat the test image as the response. Under the Lambertian assumption that each image is represented as a linear combination of other images of its subject in the dictionary, the problem of face recognition can be addressed by finding the block, i.e., the subject, whose feature matches the test image, or equivalently, by reconstructing a signal with block sparsity one.
We evaluate and compare the performances of GPSP with other algorithms using the Extended Yale B dataset [LHK05] derived from [GBK01]. This dataset consists of face images belonging to 38 human subjects. Each subject is represented by 64 face images, which are captured under various illuminating conditions. The first row of Figure 2 shows 8 images of a specific subject; they exemplify the variations in the dataset, including contrast change (the 2nd image), minor differences in facial expressions (the 3rd image), and extremely low lighting (the 5th image). For each subject, we randomly sampled $M$ images to form our training dataset of size $38\times M$, and we use the rest images for evaluation. Upon stretching each image to a vector of length $32256\,(=192\times 168)$, we employ three strategies to reduce the dimension to $D=132$ as in [EV12]:
@@
Table 2: Comparison study on the task of face recoginition using the Extended Yale B dataset [LHK05]. The first row shows sample images of a single subject. The table records the average values of frequency of exact recovery over 10 trials for different training size per subject $M=9,18,25,32$ and three dimension reduction strategies: principle component analysis (PCA), random projection (randomProj), and down-sampling (downSample). The highest levels of accuracy are bolded, and the second best are underlined.
@@ 5.5 Application to PDE Identification
GPSP was first proposed in [HKL+22] for identifying PDEs with time-space dependent coefficients from trajectory data. It was shown that GPSP can find the underlying models for various important classes of PDEs with higher success rates than BSP. In this experiment, we extend the comparison by including also BOMP, BOMPR, and BCoSaMP. We consider the viscous Burgers’ equation
@@
with $\phi(t)=0.5+0.5\tanh(10(t-0.15))$. Following the numerical settings in [HKL+23], we discretize the computational domain with spatial step 0.02 and temporal step $3\times 10^{-4}$ with which the PDE is solved. Then we downsample the data to spatial step 0.04 and temporal step $3\times 10^{-3}$, such that the data is over a $50\times 100$ grid.
The PDE (5.1) has non-trivial behaviors. The coefficient $a(x,t)$ is visualized in Figure 7 (a), and the corresponding solution obtained by the spectral method (See [HKL+22]) is shown in (b). It models a non-linear advection with a homogeneous diffusion on a circle. The speed of the advection is affected by the solution itself as well as the varying coefficient $a$. When the solution $u$ is positive, it moves from right to left; when it is negative, it moves from left to right. Moreover, the higher the magnitude of the solution is, the faster it moves locally. Consequently, solutions can form shocks without diffusion. For more details about the properties of the PDE, we refer the readers to [Eva22].
We set up a dictionary containing candidate differential operators as features to identify the PDE from the solution data. We include partial derivatives of $u$ up to order 4 and products of no more than 3 terms, leading to 55 features in the dictionary. Since the coefficients may vary both in space and time, we approximate them using linear combinations of B-spline bases of order $2$. As a result, each feature is represented by a block matrix whose columns are associated with local bases. We use 7 bases for space and 8 bases for time, leading to a block size of 56. The details of the implementation are explained in [HKL+22]. Assuming that we know the sparsity is 2, we apply BOMP, BOMPR, BCoSaMP, BSP, and GPSP to find the features. Table 3 shows the identified features by these methods. In this experiment, only GPSP finds the correct features, giving the underlying viscous Burgers’ equation. BOMP and BOMPR have identical results; they find the correct second-order differential operator, yet the non-linear part is wrong. BCoSaMP and BSP fail to identify any correct features, and the respective models are more complicated than the true equation. This comparison verifies the effectiveness of GPSP in the application of PDE identification.